# Improved 3D Scene Stylization via Text-Guided Generative Image Editing with Region-Based Control

### v1

@@ 3.3.1 Importance-Weighted Sliced Wasserstein Distance Loss
Based on the assumption that a projection direction that gives a higher 1D Wasserstein distance value is more informative, we propose applying the energy-based SWD variant with importance sampling as studied by [NH23]. Specifically, we find that extending the loss function in 15 with importance weighting using Softmax function leads to enhanced training efficiency only with a minor implementation cost compared to the vanilla SWD loss:
@@ 3.3.2 Multi-Region Importance-Weighted Sliced Wasserstein Distance Loss
Additionally, we propose a region-based style transfer loss, which is similar to the concept introduced by a 2D neural texture synthesis method [LXY22] and formulate the idea as follows.
@@ 3.3.3 Content Loss
To maintain the original structure of the scene, we further incorporate a content preservation loss [LXY22] by calculating the mean-squared error of VGG19 activations:
@@ 4.1 Experimental Setup
We evaluate our method on multiple scenes selected from Instruct-NeRF2NeRF dataset [HTE23] and Mip-NeRF360 dataset [BMV22], which provide diverse and high-quality 3D scenes. Our 2D multi-view stylization pipeline is built on top of an SDXL-based style-aligned diffusion framework [HVFCO24], enabling text-driven appearance editing. To obtain the depth maps for the tiled control image, we use MiDAS [RLH20]. For 3D scene refinement, we clone the official implementation of 2D Gaussian Splatting (2DGS) [HYC24] as our representation backbone. We use SAM2 [RGH24] to obtain segmentation masks for our multi-region loss in the stylization process. All baseline methods are evaluated using the publicly released code provided by their respective authors. We train all scenes on a single NVIDIA A100 GPU, running for 500–1000 iterations, which takes approximately 10-15 minutes depending on the image resolution and number of views in the scene.
@@ 4.2 Text-Driven 3D Stylization
We compare our method against two state-of-the-art approaches: Style-NeRF2NeRF [FMH24] and DGE [CLV24]. Both methods follow a similar approach, where the source 3D scene is directly edited using a set of consistently stylized multi-view images generated from text prompts. While sharing this overall structure, our method introduces key improvements in the multi-view generation pipeline and 3D refinement stage, leading to improved visual fidelity and semantic alignment with the input prompts. Qualitative comparisons and quantitative results are summarized in Table 1 and Figure 4.
@@ 4.2.1 Qualitative Results
We present qualitative comparisons of our 3D stylization results in Figure 4, alongside results from prior methods. Since we use the same diffusion backbone (SDXL [PEL23]) as Style-NeRF2NeRF [FMH24] and fix the random seed across experiments, the overall stylization atmosphere appears similar. However, our method yields results that are more faithful to the style prompts, with sharper details and fewer visual artifacts. While DGE demonstrates strong performance in portrait-style edits, it occasionally fails to stylize the full scene coherently and often produces over-saturated colors. For a more comprehensive view, we encourage readers to watch the demo video.
@@ 4.2.2 Quantitative Results
We quantitatively evaluate the performance of our method using standard metrics following prior work [GPM22], as summarized in Table 1. Specifically, we report CLIP similarity, which measures image-text alignment between the source and stylized images in CLIP’s embedding space. To assess view consistency, we also compute warping errors across sequential frames using optical flow estimated by RAFT [TD20], which captures temporal coherence from a geometric perspective. We calculate the average mean squared error (MSE) between the forward-warped i1i-1-th frame and the $i$-th frame. As shown in Table 1, our method achieves competitive performance across all metrics when compared to baseline approaches. That said, we acknowledge that stylization is inherently subjective, and user preferences may vary depending on specific goals. To complement the quantitative evaluation, we conducted a user study involving 58 participants. Each participant was shown 12 randomly ordered scene–style prompt pairs (4 scenes × 3 styles), and asked to choose the most visually appealing result for each. In total, 696 responses were collected (58 participants × 12 questions), and the aggregated preferences are also reported in Table 1.
Method
CLIP
Similarity $\uparrow$
Warp
Error $\downarrow$
User
Pref. $\uparrow$
Style-N2N
w/ GS [FMH24]
0.142 0.054 19.9%
DGE
[CLV24]
0.184 0.072 22.0%
@@
Table 1: Method Comparisons. We report quantitative values using metrics following common literature. For the Style-NeRF2NeRF baseline [FMH24], we replace their backbone 3D representation with the same 2D Gaussian splatting [HYC24] as ours to make a fair comparison.
@@ 4.3 Ablations
To validate the effectiveness of our proposed components, we perform ablation experiments by comparing our full method with several simplified variants. The first variant replaces our tiled depth map reference with a single depth map and employs a fully-shared attention diffusion pipeline, as proposed in prior work [FMH24]. The second variant turns off the multi-region functionality by omitting the input masks, allowing us to assess the contribution of our multi-region SWD loss. We also verify that skipping content loss substantially increases the Warp Error, proving that content loss contributes to maintaining 3D view consistency.
@@
Variant
CLIP
Similarity $\uparrow$
Warp
Error $\downarrow$
w/o tiled ref. MV pipeline
0.178 0.061
w/o multi-region loss
0.201 0.058
w/o content loss
0.210 0.064
@@ Acknowledgements
Figure 4: Method Comparison. We compare our method against Style-NeRF2NeRF [FMH24] and DGE [CLV24]. As shown, our approach produces clearer and more visually artistic results that are faithful to the given style prompts, while exhibiting fewer artifacts. For a fair comparison, we replace the underlying NeRF representation in the Style-NeRF2NeRF baseline with the same 2D Gaussian Splatting (2DGS) [HYC24] used in our method.
@@
Figure 7: Example of 3D scene stylization using our multi-region loss. Three distinct styles are applied to different semantic regions based on the segmentation masks, enabling spatially controlled stylization. In this example, the background region is excluded from stylization and remains unchanged, demonstrating the ability to selectively preserve parts of the original scene.
@@ A Additional Implementation Details
We implemented our 2D multi-view stylization pipeline by building on the code from [HVFCO24], which uses Stable Diffusion XL [PEL23] as a backbone. We attach depth-conditioning ControlNet [ZRA23] released by Diffusers from Huggingface. For view generation, we set the ControlNet conditioning scale to $1.0$ with $50$ diffusion steps. We also find that a classifier guidance scale within the range $[4.5,13.5]$ usually works well, depending on the stylization objective. We encourage users to test our released code for further details.
@@ A.2 Loss Functions
For pre-training of the source scene, we follow the same setup as [HYC24]. During the scene stylization stage, we disable the L1 photometric loss and add our proposed IWD-SWD loss, which calculates the (importance-weighted) sliced Wasserstein distance between VGG19 feature maps of the rendered image and the corresponding ground truth image. We use the first 12 layers of VGG19 to extract the feature maps. The number of projections are 3, 6, 13, and 26 ($5\%$ of 64, 128, 256, and 512) for layer $l=1,2$, $l=3,4$, $l=5,6,7,8$, and $l=9,10,11,12$ respectively. The code for IW-SWD is based on the implementation released by [HVCB21].
@@ B Importance-Weighted Sliced Wasserstein Distance
In this section, we leave supplementary notes regarding the formulation of the Importance-Weighted Sliced Wasserstein Distance (IW-SWD). For readers interested in more theoretical details, we strongly recommend referring to the original paper [NH23].
@@ B.2 Energy-Based Sliced Wasserstein
Although $\widehat{SW}(\mu,\nu)$ serves as a useful term for general machine learning tasks, some slicing directions from uniform sampling may not be as informative towards the true expectation value $SW(\mu,\nu)$. To address this issue, [NH23] introduced the Energy-Based Sliced Wasserstein (EBSW) distance as an optimization-free variant that exploits the value of the projected Wasserstein distance between $\mu$ and $\nu$.
@@
The authors also propose several sampling schemes for EBSW approximation. Among them, we use the Monte Carlo sampling based on importance sampling [KVD78]. The EBSW distance using a sampling proposal distribution $\sigma_{0}(V)\in\mathcal{P}(\mathbb{S}^{d-1})$ can be reformulated as the following:
@@
Specifically, IW-SWD corresponds to this IS-EBSW case presented in the paper [NH23], where we adopt the exponential energy function $f(x)=e^{x}$ with a uniform proposal distribution $\sigma_{0}(V)=\mathcal{U}(\mathbb{S}^{d-1})$, which is equivalent to weighting the SWD values using the Softmax function.
@@ C Additional Comparisons
In figure 8 and 9, we include additional visual comparisons against Instruct-GS2GS [VH24], GaussianEditor [CCZ24], VcEdit [WYW24], and Style-NeRF2NeRF (w/ original NeRF representation as baseline) [FMH24] for reference. The results are produced using the default settings for each method. Therefore, please note that tuning parameters such as the guidance scale of the underlying Instruct-Pix2Pix model [BHE23] may result in different outcomes. We recommend that users test different methods according to their varying goals and requirements.
Figure 8: Additional Method Comparison 1. Our method exhibits artistically superior results without blurry artifacts. While VcEdit [WYW24] can perform convincing editing in some cases, we found it to be memory-intensive ( 80GB of GPU memory with full training views) and relatively difficult for large-scale 360-degree scenes to converge when using holistic style transfer prompts.

### v2

@@ 3.3.1 Importance-Weighted Sliced Wasserstein Distance Loss
Based on the assumption that a projection direction that gives a higher 1D Wasserstein distance value is more informative, we propose applying the energy-based SWD variant with importance sampling as studied by [nguyen2023energy]. Specifically, we find that extending the loss function in 15 with importance weighting using Softmax function leads to enhanced training efficiency only with a minor implementation cost compared to the vanilla SWD loss:
@@ 3.3.2 Multi-Region Importance-Weighted Sliced Wasserstein Distance Loss
Additionally, we propose a region-based style transfer loss, which is similar to the concept introduced by a 2D neural texture synthesis method [li2022sliced] and formulate the idea as follows.
@@ 3.3.3 Content Loss
To maintain the original structure of the scene, we further incorporate a content preservation loss [li2022sliced] by calculating the mean-squared error of VGG19 activations:
@@ 4.1 Experimental Setup
We evaluate our method on multiple scenes selected from Instruct-NeRF2NeRF dataset [instructnerf2023] and Mip-NeRF360 dataset [barron2022mip], which provide diverse and high-quality 3D scenes. Our 2D multi-view stylization pipeline is built on top of an SDXL-based style-aligned diffusion framework [hertz2024style], enabling text-driven appearance editing. To obtain the depth maps for the tiled control image, we use MiDAS [ranftl2020towards]. For 3D scene refinement, we clone the official implementation of 2D Gaussian Splatting (2DGS) [huang20242d] as our representation backbone. We use SAM2 [ravi2024sam2] to obtain segmentation masks for our multi-region loss in the stylization process. All baseline methods are evaluated using the publicly released code provided by their respective authors. We train all scenes on a single NVIDIA A100 GPU, running for 500–1000 iterations, which takes approximately 10-15 minutes depending on the image resolution and number of views in the scene.
@@ 4.2 Text-Driven 3D Stylization
We compare our method against two state-of-the-art approaches: Style-NeRF2NeRF [fujiwara2024sn2n] and DGE [chen2024dge]. Both methods follow a similar approach, where the source 3D scene is directly edited using a set of consistently stylized multi-view images generated from text prompts. While sharing this overall structure, our method introduces key improvements in the multi-view generation pipeline and 3D refinement stage, leading to improved visual fidelity and semantic alignment with the input prompts. Qualitative comparisons and quantitative results are summarized in Table 1 and Figure 4.
@@ 4.2.1 Qualitative Results
We present qualitative comparisons of our 3D stylization results in Figure 4, alongside results from prior methods. Since we use the same diffusion backbone (SDXL [podell2023sdxl]) as Style-NeRF2NeRF [fujiwara2024sn2n] and fix the random seed across experiments, the overall stylization atmosphere appears similar. However, our method yields results that are more faithful to the style prompts, with sharper details and fewer visual artifacts. While DGE demonstrates strong performance in portrait-style edits, it occasionally fails to stylize the full scene coherently and often produces over-saturated colors. For a more comprehensive view, we encourage readers to watch the demo video.
@@ 4.2.2 Quantitative Results
We quantitatively evaluate the performance of our method using standard metrics following prior work [gal2022stylegan], as summarized in Table 1. Specifically, we report CLIP similarity, which measures image-text alignment between the source and stylized images in CLIP’s embedding space. To assess view consistency, we also compute warping errors across sequential frames using optical flow estimated by RAFT [teed2020raft], which captures temporal coherence from a geometric perspective. We calculate the average mean squared error (MSE) between the forward-warped i1i-1-th frame and the $i$-th frame. As shown in Table 1, our method achieves competitive performance across all metrics when compared to baseline approaches. That said, we acknowledge that stylization is inherently subjective, and user preferences may vary depending on specific goals. To complement the quantitative evaluation, we conducted a user study involving 58 participants. Each participant was shown 12 randomly ordered scene–style prompt pairs (4 scenes × 3 styles), and asked to choose the most visually appealing result for each. In total, 696 responses were collected (58 participants × 12 questions), and the aggregated preferences are also reported in Table 1.
Method CLIP Similarity $\uparrow$ Warp Error $\downarrow$ User Pref. $\uparrow$
Style-N2N w/ GS [fujiwara2024sn2n] 0.142 0.054 19.9%
DGE [chen2024dge] 0.184 0.072 22.0%
@@
Table 1: Method Comparisons. We report quantitative values using metrics following common literature. For the Style-NeRF2NeRF baseline [fujiwara2024sn2n], we replace their backbone 3D representation with the same 2D Gaussian splatting [huang20242d] as ours to make a fair comparison.
@@ 4.3 Ablations
To validate the effectiveness of our proposed components, we perform ablation experiments by comparing our full method with several simplified variants. The first variant replaces our tiled depth map reference with a single depth map and employs a fully-shared attention diffusion pipeline, as proposed in prior work [fujiwara2024sn2n]. The second variant turns off the multi-region functionality by omitting the input masks, allowing us to assess the contribution of our multi-region SWD loss. We also verify that skipping content loss substantially increases the Warp Error, proving that content loss contributes to maintaining 3D view consistency.
@@
Variant CLIP Similarity $\uparrow$ Warp Error $\downarrow$
w/o tiled ref. MV pipeline 0.178 0.061
w/o multi-region loss 0.201 0.058
w/o content loss 0.210 0.064
@@ Acknowledgements
Figure 4: Method Comparison. We compare our method against Style-NeRF2NeRF [fujiwara2024sn2n] and DGE [chen2024dge]. As shown, our approach produces clearer and more visually artistic results that are faithful to the given style prompts, while exhibiting fewer artifacts. For a fair comparison, we replace the underlying NeRF representation in the Style-NeRF2NeRF baseline with the same 2D Gaussian Splatting (2DGS) [huang20242d] used in our method.
@@
Figure 7: Example of 3D scene stylization using our multi-region loss. Three distinct styles are applied to different semantic regions based on the segmentation masks, enabling spatially controlled stylization. In this example, the background region is excluded from stylizatixon and remains unchanged, demonstrating the ability to selectively preserve parts of the original scene.
@@ A Additional Implementation Details
We implemented our 2D multi-view stylization pipeline by building on the code from [hertz2024style], which uses Stable Diffusion XL [podell2023sdxl] as a backbone. We attach depth-conditioning ControlNet [zhang2023adding] released by Diffusers from Huggingface. For view generation, we set the ControlNet conditioning scale to $1.0$ with $50$ diffusion steps. We also find that a classifier guidance scale within the range $[4.5,13.5]$ usually works well, depending on the stylization objective. We encourage users to test our released code for further details.
@@ A.2 Loss Functions
For pre-training of the source scene, we follow the same setup as [huang20242d]. During the scene stylization stage, we disable the L1 photometric loss and add our proposed IWD-SWD loss, which calculates the (importance-weighted) sliced Wasserstein distance between VGG19 feature maps of the rendered image and the corresponding ground truth image. We use the first 12 layers of VGG19 to extract the feature maps. The number of projections are 3, 6, 13, and 26 ($5\%$ of 64, 128, 256, and 512) for layer $l=1,2$, $l=3,4$, $l=5,6,7,8$, and $l=9,10,11,12$ respectively. The code for IW-SWD is based on the implementation released by [heitz2021sliced].
@@ B Importance-Weighted Sliced Wasserstein Distance
In this section, we leave supplementary notes regarding the formulation of the Importance-Weighted Sliced Wasserstein Distance (IW-SWD). For readers interested in more theoretical details, we strongly recommend referring to the original paper [nguyen2023energy].
@@ B.2 Energy-Based Sliced Wasserstein
Although $\widehat{SW}(\mu,\nu)$ serves as a useful term for general machine learning tasks, some slicing directions from uniform sampling may not be as informative towards the true expectation value $SW(\mu,\nu)$. To address this issue, [nguyen2023energy] introduced the Energy-Based Sliced Wasserstein (EBSW) distance as an optimization-free variant that exploits the value of the projected Wasserstein distance between $\mu$ and $\nu$.
@@
The authors also propose several sampling schemes for EBSW approximation. Among them, we use the Monte Carlo sampling based on importance sampling [kloek1978bayesian]. The EBSW distance using a sampling proposal distribution $\sigma_{0}(V)\in\mathcal{P}(\mathbb{S}^{d-1})$ can be reformulated as the following:
@@
Specifically, IW-SWD corresponds to this IS-EBSW case presented in the paper [nguyen2023energy], where we adopt the exponential energy function $f(x)=e^{x}$ with a uniform proposal distribution $\sigma_{0}(V)=\mathcal{U}(\mathbb{S}^{d-1})$, which is equivalent to weighting the SWD values using the Softmax function.
@@ C Additional Comparisons
In figure 8 and 9, we include additional visual comparisons against Instruct-GS2GS [igs2gs], GaussianEditor [chen2024gaussianeditor], VcEdit [wang2024view], and Style-NeRF2NeRF (w/ original NeRF representation as baseline) [fujiwara2024sn2n] for reference. The results are produced using the default settings for each method. Therefore, please note that tuning parameters such as the guidance scale of the underlying Instruct-Pix2Pix model [brooks2023instructpix2pix] may result in different outcomes. We recommend that users test different methods according to their varying goals and requirements.
Figure 8: Additional Method Comparison 1. Our method exhibits artistically superior results without blurry artifacts. While VcEdit [wang2024view] can perform convincing editing in some cases, we found it to be memory-intensive ( 80GB of GPU memory with full training views) and relatively difficult for large-scale 360-degree scenes to converge when using holistic style transfer prompts.